# Shear banding as a cause of non-monotonic stress relaxation after flow cessation

### v1

@@ Abstract
To date, such observations remain poorly understood theoretically. Indeed, it is difficult to intuit a mechanism via which stress can increase over any time interval, once shear driving no longer acts on the material. In associating gels [21], it was suggested to arise via the reformation of broken bonds after flow cessation. The released energy was argued to then increase the system’s elastic energy via the formation of domains partly aligned with the prior shear. Within a phenomenological structural kinetic model [25], it was suggested to arise via the competition of viscoelastic aging with rejuvenation driven by the viscous rate of strain. In the context of constitutive modeling of thixotropy, the buildup of a material’s modulus over time has been argued to violate the second law of thermodynamics [26], although elasticity was argued to contribute significantly to thixotropy in Ref. [27].
In this Letter, we put forward an alternative explanation for the phenomenon of non-monotonic stress relaxation after flow cessation, arguing that the widespread physical ingredients of (i) coexisting shear bands of differing viscosity prior to flow cessation and (ii) frustrated local stresses can interact to give rise to it. We substantiate this scenario by simulating the widely used soft glassy rheology model, in a form extended to allow shear banding. We furthermore offer a detailed mechanistic understanding of the basic underlying physics, which can be understood via the evolving distributions of frustrated local stresses in the two former bands, governed by their differing effective temperatures.
@@
Results — We start in Fig. 1a) by showing the flow curves of shear stress as a function of shear rate in steady shear prior to switch-off, $\sigma_{\rm ss}(\bar{\dot{\gamma}})$. Solid lines correspond to systems in which the shear field is allowed to become heterogeneous. Dashed lines show the underlying constitutive curves, with the shear artificially constrained to be uniform. For $a>a^{*}\approx 0.84$, the constitutive curves are non-monotonic, with a regime of negative slope $d\sigma/d\dot{\gamma}<0$ that causes shear bands to form over a window of shear rates $0<\bar{\dot{\gamma}}<\bar{\dot{\gamma}}_{\rm c}(a)$. Over this window, the flow curves are flat and differ from the homogeneous constitutive curves. For $\bar{\dot{\gamma}}>\bar{\dot{\gamma}}_{\rm c}$, the shear is again homogeneous, and the constitutive and flow curves coincide. In principle, the flat stress plateau should extend rightwards until it intersects the constitutive curve. In practice, for higher shear rates in this regime the system gets stuck in a state of metastable homogeneous shear, giving a slight step in the flow curve.
@@
For one particular imposed shear rate, $\dot{\gamma}=0.16$, shown by dots in Fig. 1a, we also explore the effect of varying the delay time $\tau_{x}$ (Fig. 3a) and the coupling parameter $a$ (Fig. 3b). As can be seen, to obtain a non-monotonic stress decay after flow cessation, $\nu>0$, a sufficiently large value of the coupling parameter is required, $a\geq a^{**}>a^{*}$. This ensures that the flow state prior to switch-off is shear banded, and with a sufficiently large effective temperature difference between the bands. For $a>a^{**}\approx 1.5$, $\nu$ then increases with increasing $a$. Also required is a sufficiently long delay time for the relaxation of the effective temperature post switch-off, $\tau_{x}\geq\tau_{x}^{*}\approx 5.0$. For $\tau_{x}>\tau_{x}^{*}$, $\nu$ then increases with $\tau_{x}$.
@@
Figure 4: a) Non-monotonic stress decay as a function of time $t$ after flow cessation. b) Distribution of local strains/stresses in the solid-like region, while the macroscopic stress is decreasing, denoted by crosses in a). The distribution shifts elastically leftwards to lower local strain values. c) Corresponding local strain distribution while the macroscopic stress is increasing, denoted by squares in a). Here elements with negative values of $l$ plastically yield, resulting in the increase in macroscopic stress. Parameter values are $a=6,\tau_{x}=100,\bar{\dot{\gamma}}=0.08$ and $\lambda=0.005$.
Mechanistic understanding — We now explain the physical mechanism of this phenomenon with reference to the distribution of local strains $l$, or equivalently of local stresses $kl$ (recall $k=1$.) This is plotted in Fig. 4 at several times after flow cessation. Consider a shear banded state before switch-off, with averages over elements in the high shear band and unsheared band denoted $\langle\cdot\rangle_{\rm h}$ and $\langle\cdot\rangle_{\rm u}$. In the high shear band, the strong shear will have fluidised the material, with elements occupying shallow traps of high average yielding rate $\langle r\rangle_{\rm h}$. In the unsheared band the material will be solid-like, elements having aged into deep traps with low $\langle r\rangle_{\rm u}$. Force balance however demands a uniform shear stress, $\sigma=\langle l\rangle_{\rm h}=\langle l\rangle_{\rm u}$.
@@
In the fluidised region, the rate of plastic stress relaxation $\langle lr\rangle_{\rm h}$ will be high. This results in a stress decay, giving negative $\dot{\sigma}$. Force balance demands the same negative $\dot{\sigma}$ in the solid-like region, in which $\langle lr\rangle_{\rm u}$ is small. A stress decay (negative $\dot{\sigma}$) thus demands a negative $\dot{\gamma}_{\rm u}$, i.e., a backwards elastic recoil of the solid-like region, as the material unloads. This elastic recoil advects the local strains of all elements in that solid-like region to lower values (Fig. 4b) – including into the negative range of $l$. Over a subsequent, slower timescale, these elements then plastically yield (Fig. 4c). In particular, those elements that yield from negative $l$ values and reset their $l$ values to zero contribute a negative rate of stress relaxation $\langle lr\rangle_{\rm u}$, giving a positive contribution to the stress rate of change, $\dot{\sigma}=\dot{\gamma}_{\rm u}-\langle lr\rangle_{\rm u}$. It is this plastic yielding of elements with negative $l$ values that can give rise to the observed increase in the macroscopic stress. It does so for systems in which the shear bands prior to flow cessation had a sufficiently large difference in effective temperature, and in which the decay time for relaxation after flow cessation is sufficiently slow. This results in a separation of timescales between (first) the initially fast plastic yielding of positive local stresses from the fluidised band, which causes the initial rapid decay of the macroscopic stress; and (second) the later plastic yielding of negative local stresses from the solid band, which causes the slow rise in macroscopic stress.

### v2

@@ Abstract
To date, such observations remain poorly understood theoretically. Indeed, it is difficult to intuit a mechanism via which stress can increase over any time interval, once shear driving no longer acts on the material. In associating gels [21], it was suggested to arise via the reformation of broken bonds after flow cessation. The released energy was argued to then increase the system’s elastic energy via the formation of domains partly aligned with the prior shear. Within a phenomenological structural kinetic model [25], the stress increase was predicted to arise due to the competition of viscoelastic aging with rejuvenation. In the context of constitutive modeling of thixotropy, the buildup of a material’s modulus over time has been argued to violate the second law of thermodynamics [26], although elasticity was argued to contribute significantly to thixotropy in Ref. [27].
In this Letter, we put forward an alternative explanation for the phenomenon of non-monotonic stress relaxation after flow cessation. In particular, we argue that the widespread physical ingredients of (i) coexisting shear bands of differing viscosity prior to flow cessation and (ii) frustrated local stresses can interact to give rise to it. We substantiate this scenario by simulating the widely used soft glassy rheology model, in a form extended to allow shear banding. It is known that shear banding is present in a number of the systems that exhibit the stress increases [28]. We furthermore offer a detailed mechanistic understanding of the basic underlying physics, which can be understood via the evolving distributions of frustrated local stresses in the two former bands, governed by their differing effective temperatures.
@@
Results — We start in Fig. 1a) by showing the flow curves of shear stress as a function of shear rate in steady shear prior to switch-off, $\sigma_{\rm ss}(\bar{\dot{\gamma}})$. Solid lines correspond to systems in which the shear field is allowed to become heterogeneous. Dashed lines show the underlying constitutive curves, with the shear artificially constrained to be uniform. For $a>a^{*}\approx 0.84$, the constitutive curves are non-monotonic, with a regime of negative slope $d\sigma/d\dot{\gamma}<0$ that causes shear bands to form over a window of shear rates $0<\bar{\dot{\gamma}}<\bar{\dot{\gamma}}_{\rm c}(a)$. Over this window, the flow curves are flat and differ from the homogeneous constitutive curves. The strain rates in the two shear bands that form in this regime correspond to the shear rates at each end of the plateau in the steady state flow curve. The proportion of the gap occupied by each band is such that the shear rate averaged over the two bands, across the gap, equals the applied strain rate. The interface between the two bands has a slightly diffuse width governed by the prefactor $\lambda$ in the final term of Eqn. 1. For $\bar{\dot{\gamma}}>\bar{\dot{\gamma}}_{\rm c}$, the shear is again homogeneous, and the constitutive and flow curves coincide. In principle, the flat stress plateau should extend rightwards until it intersects the constitutive curve. In practice, for higher shear rates in this regime the system gets stuck in a state of metastable homogeneous shear, giving a slight step in the flow curve.
@@
For one particular imposed shear rate, $\dot{\gamma}=0.16$, shown by dots in Fig. 1a, we also explore the effect of varying the delay time $\tau_{x}$ (Fig. 3a) and the coupling parameter $a$ (Fig. 3b). As can be seen, to obtain a non-monotonic stress decay after flow cessation, $\nu>0$, a sufficiently large value of the coupling parameter is required, $a\geq a^{**}>a^{*}$. This ensures that the flow state prior to switch-off is shear banded, and with a sufficiently large effective temperature difference between the bands. For $a>a^{**}\approx 1.5$, $\nu$ then increases with increasing $a$. Also required is a sufficiently long delay time for the relaxation of the effective temperature post switch-off, $\tau_{x}\geq\tau_{x}^{*}\approx 5.0$. For $\tau_{x}>\tau_{x}^{*}$, $\nu$ then increases with $\tau_{x}$. The time at which the stress reaches a minimum is largely independent of the imposed shear rate and the coupling parameter $a$, but increases linearly with the parameter $\tau_{x}$ that governs the timescale for the evolution of the effective temperature in Eqn. 1.
@@
[…]
Figure 4: a) Non-monotonic stress decay as a function of time $t$ after flow cessation. The distributions of local stresses are shown in panels b+d) at times denoted by crosses during the initial stress decrease in a). Panels c+e) show the distributions during the later stress increase, at times denoted by squares in a). Panels b+c) pertain to the region previously (during flow) occupied by the high shear band. Panels (d+e) pertain to the region previously occupied by the low shear band. b) In the liquid-like region, elements with positive values of $l$ plastically yield at early times, causing the stress to relax. c) At longer times in this region, the distribution shifts slightly to higher local strain values. d) In the solid-like region, the distribution shifts elastically leftwards to lower local strain values during the initial stress relaxation. e) At longer times in the solid-like region, elements with negative values of $l$ plastically yield. It is this yielding of negative local stresses that results in the increase in macroscopic stress. Parameter values are $a=6,\tau_{x}=100,\bar{\dot{\gamma}}=0.08$ and $\lambda=0.005$.
Mechanistic understanding — We now explain the physical mechanism of this phenomenon. We do so with reference to the distribution of local strains $l$, or equivalently of local stresses $kl$ (recall $k=1$). This is plotted at several times after flow cessation in the region previously (during flow) occupied by the high shear band in Figs. 4b,c); and in the region previously occupied by the low shear band in Figs. 4d,e). We have checked that the width of these two regions stays relatively constant during the stress relaxation post-cessation.
Consider a shear banded state before switch-off, with averages over elements in the high shear band and unsheared band denoted $\langle\cdot\rangle_{\rm h}$ and $\langle\cdot\rangle_{\rm u}$. In the high shear band, the strong shear will have fluidised the material, with elements occupying shallow traps of high average yielding rate $\langle r\rangle_{\rm h}$. In the unsheared band the material will be solid-like, elements having aged into deep traps with low $\langle r\rangle_{\rm u}$. Force balance however demands a uniform shear stress, $\sigma=\langle l\rangle_{\rm h}=\langle l\rangle_{\rm u}$.
@@
In the fluidised region, the rate of plastic stress relaxation $\langle lr\rangle_{\rm h}$ will be high. Accordingly, the initially broad distribution of values of $l$ (purple line) quickly relaxes to a much narrower distribution focused nearer the origin (red line), as shown in Fig. 4b). This results in a rapid stress decay, giving negative $\dot{\sigma}$. Force balance demands the same negative $\dot{\sigma}$ in the solid-like region, in which $\langle lr\rangle_{\rm u}$ is small. A stress decay (negative $\dot{\sigma}$) thus demands a negative $\dot{\gamma}_{\rm u}$, i.e., a backwards elastic recoil of the solid-like region, as the material unloads. This elastic recoil advects the local strains of all elements in that solid-like region to lower values (Fig. 4d) – including into the negative range of $l$. Over a subsequent, slower timescale, these elements then plastically yield (Fig. 4e). In particular, those elements that yield from negative $l$ values and reset their $l$ values to zero contribute a negative rate of stress relaxation $\langle lr\rangle_{\rm u}$, giving a positive contribution to the stress rate of change, $\dot{\sigma}=\dot{\gamma}_{\rm u}-\langle lr\rangle_{\rm u}$. It is this plastic yielding of elements with negative $l$ values that can give rise to the observed increase in the macroscopic stress. It does so for systems in which the shear bands prior to flow cessation had a sufficiently large difference in effective temperature, and in which the decay time for relaxation after flow cessation is sufficiently slow. This results in a separation of timescales between (first) the initially fast plastic yielding of positive local stresses from the fluidised band, which causes the initial rapid decay of the macroscopic stress; and (second) the later plastic yielding of negative local stresses from the solid band, which causes the slow rise in macroscopic stress.